# Embedding Ontologies via Incorporating Extensional and Intensional Knowledge

### v4

@@ 5.3 Triple Classification
Table 3: Experimental results of SubClassOf triple classification(%). Results of * are taken from [43]
@@
Table 4: Experimental results on link prediction and triple classification for relational triples on YAGO39K and DB99K242. Hits@N uses results of “Filter” evaluation setting. Results of * are taken from [43]
@@ 5.4 Link Prediction
Link prediction aims to predict the missing head entity (?, r, t) or tail entity (h, r, ?). Same as TransC [20] and [43], We only tested link prediction on Relational triples because the InstanceOf relation and SubClassOf relation are not suitable for this task. Following most previous works [20] [2] [41], we adopt two evaluation metrics as followings: (1) the mean reciprocal rank of all correct instances (MRR); (2) the proportion of correct instances that rank no larger than N (Hits@N). Since a corrupted triple may have already existed in knowledge base, which should be regarded as a correct prediction. The settings “Raw” and “Filter” distinguish whether or not to consider the impact of a corrupted triple already existing in the knowledge base.
@@ 6 Conclusion and Future Work
Author Contributions
K.Y. Wang(ky-wang@seu.edu.cn), G.L. Qi (gqi@seu.edu.cn) and J.Y. Chen were responsible for the design of the research, the implementation of the approach, collection and analysis the data. K.Y. Wang(ky-wang@seu.edu.cn) was responsible for the writing of the manuscript. Huang( huangyi@chinamobile.com) and T.X. Wu(tianxingwu@seu.edu.cn) were responsible for the revision of the manuscript. All the authors have made meaningful and valuable contributions in revising and proofreading the resulting manuscript.
Acknowledgements
Research in this paper was partially supported by the Natural Science Foundation of China (U21A20488). We thank the Big Data Computing Center of Southeast University for providing the facility support on the numerical calculations in this paper.

### v7

@@ 5.1 Datasets
Table 3: Experimental results of SubClassOf triple classification(%). Results of * are taken from Yu et al. (2023)
@@
Table 4: Experimental results on link prediction and triple classification for relational triples on YAGO39K and DB99K242. Hits@N uses results of “Filter” evaluation setting. Results of * are taken from Yu et al. (2023)
5.2 Experimental Settings
Followed by Yu et al. (2023), we selected 13 state-of-the-art knowledge representation learning models as baselines for evaluating our proposed approach. These models are primarily categorized into two classes: (1) Methods distinguishing concepts and instances, including TransC Lv et al. (2018), TransFG Yu et al. (2019), JECI Wang et al. (2020), TransEllipsoid Yu et al. (2023) and TransCuboid Yu et al. (2023), (2) methods that do not distinguish between concepts and instances, such as TransE Bordes et al. (2013), TransH Wang et al. (2014), TransR Lin et al. (2015), TransD Ji et al. (2015), DistMult Yang et al. (2014), HolE Nickel et al. (2016), ComplEx Trouillon et al. (2016) and Concept2Vec Qiu and Wang (2020).
All data and codes can be found in supplementary material.
5.3 Triple Classification
@@ 5.4 Link Prediction
Link prediction aims to predict the missing head entity (?, r, t) or tail entity (h, r, ?). Same as TransC Lv et al. (2018) and Yu et al. (2023), We only tested link prediction on Relational triples because the InstanceOf relation and SubClassOf relation are not suitable for this task. Following most previous works Lv et al. (2018) Bordes et al. (2013) Yao et al. (2019), we adopt two evaluation metrics as followings: (1) the mean reciprocal rank of all correct instances (MRR); (2) the proportion of correct instances that rank no larger than N (Hits@N). Since a corrupted triple may have already existed in knowledge base, which should be regarded as a correct prediction. The settings “Raw” and “Filter” distinguish whether or not to consider the impact of a corrupted triple already existing in the knowledge base.